# Energy Matching: Unifying Flow Matching and Energy-Based Models for Generative Modeling

### v3

@@ A.1 Energy landscape during training
In Figure 5, we visualize how the potential $V_{\theta}(x)$ transitions from a flow-like regime, where the OT loss enforces nearly zero curvature away from the data manifold (a), to an EBM-like regime, where the curvature around the new data geometry (here, two moons) is adaptively increased to approximate $\log p_{\text{data}}(x)$ (b). This two-stage design yields a well-shaped landscape that is both efficient to sample (thanks to a mostly flat potential between clusters) and accurate for density estimation near the data modes. For comparison, (c) shows an EBM trained solely with contrastive divergence, exhibiting sharper but less globally consistent basins.
@@
Figure 5: Visualization of the energy $V_{\theta}(x)$ landscapes driving the samples from eight Gaussians to two moons. See Figure 1 for the 2D perspective. (a) The OT flow loss enforces zero curvature in $V_{\theta}(x)$ along the trajectories to the target. (b) Around the 2 Moons, the curvature of $V_{\theta}(x)$ is adjusted to approximate $\log p_{\text{moons}}(x)\propto V_{\theta}(x)$ while remaining close to the pretrained landscape elsewhere. Combining these objectives yields a potential energy landscape that is both efficient for sampling and representative of the underlying target data distribution. (c) An EBM is shown for comparison, trained using contrastive divergence loss. Visible mode collapse that slows down the equilibration. Less regular landscape away from the data as it needs many simulations to explore it.
@@ A.2 Ablation on the sampling time
Here, we provide ablation studies on CIFAR-10 unconditional generation. Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, producing a stable Boltzmann distribution from which one can sample. Figure 6 illustrates the FID as a function of sampling time $\operatorname{\tau_{s}}$ for models trained under these different regimes. In the case of pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$; this occurs because, once the samples move close to the data manifold, there is no Boltzmann-like potential well to keep them from drifting away. Because the fidelity slope near the data manifold is steep with respect to sampling time, methods lacking explicit time-conditioning can easily overshoot or undershoot, significantly impacting fidelity. This behavior might explain why some models degrade in performance when made time-independent, as recently reported by Sun et al. [2025].
@@
In Figure 6 we also report results for different values of the temperature-switching parameter $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
11112222333300222244446666888810101010Sampling time $\operatorname{\tau_{s}}$ vs sampling time $\operatorname{\tau_{s}}$$\operatorname{\tau^{\ast}}=0.9,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\tau^{\ast}}=1.0,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
@@
Figure 6: CIFAR-10 unconditional generation FID vs. sampling time $\operatorname{\tau_{s}}$ when sampling from models trained under different scenarios: pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ and combined $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, with temperature regime switching parameter $\operatorname{\tau^{\ast}}\in\{0.9,1.0\}$. Lower FID indicates better generative quality. All results for Euler-Heun with $\Delta t=0.01$.
Appendix B Posterior sampling
This section provides additional algorithmic details and experimental settings for inverse problem sampling tasks presented in the main text.
B.1 Algorithm description
The detailed sampling procedure for inverse problems, optionally incorporating an interaction energy term to explicitly control the diversity among generated samples, is provided in Algorithm 3. The Langevin dynamics alternates between evaluating a composite potential $U_{\theta}(x)$, composed of a data fidelity term, a learned energy-based prior $V_{\theta}$, and, optionally, an interaction energy $W(x,x^{\prime})$. This interaction term can, for instance, encourage diversity among parallel chains. Samples are initialized independently from the Gaussian prior and iteratively refined through gradient-based updates. Specific hyperparameters, including the discretization step size $\Delta t$, total sampling time $T$, and the interaction energy strength, are detailed in Appendix D.
Algorithm 3 Sampling for inverse problems (with optional interaction energy)
1:for $m=1$ to $M$ do
2:    Initialize $x_{m}^{(0)}\sim\mathcal{N}(0,I)$triangleright Start each chain from Gaussian prior
3:end for
4:$N\leftarrow\lfloor\operatorname{\tau_{s}}/\Delta t\rfloor$triangleright Number of Langevin steps for sampling time $\operatorname{\tau_{s}}$
5:for $n=0$ to $N-1$ do
6:    for $m=1$ to $M$ do $\triangleright$ Data fidelity + prior + interaction energy
7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}-\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
8:         $\eta\sim\mathcal{N}(0,I)$triangleright Gaussian noise for Langevin step
9:         $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta$triangleright Langevin step
10:    end for
11:end for
12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples
B.2 AAV inverse design protein generation
@@
We consider protein fitness optimization for the adeno-associated virus (AAV) in the medium and hard data regime as proposed by [Kirjner et al., 2024]. The setup for the latent encoding as well as the backbone architecture for the learned prior in the latent space follows [Bogensperger et al., 2025]. Conditional sampling uses classifier guidance via learned predictor networks $g_{\phi}$ or $\tilde{g}_{\phi}$ to guide samples toward high-fitness regions. The CNN-based networks predicting fitness for AAV sequences are directly sourced from [Kirjner et al., 2024] and are trained only on the limited training data for the medium and hard setup, respectively.
@@
For training we follow Algorithm 1 and Algorithm 2. We then sample 128 sequences using Algorithm 3 (same batch size for all baselines), the hyperparameters are detailed in Appendix D. Generated sequences are evaluated for fitness using the learned oracle from Kirjner et al. [2024]. Further, they are assessed for diversity within the generated sequences and novelty relative to the training set [Jain et al., 2022]. As shown in Table 4, our method achieves state-of-the-art fitness while boosting diversity, and incorporating interaction energy in Algorithm 3 further increases diversity with only a minimal impact on fitness.
@@ Appendix C Details on LID estimation
$\mathrm{LID}(x_{\mathrm{data}})=d-\lim_{r\to 0}\frac{\log\bigl(M(r)\bigr{missing})}{\log(r)}.$
@@
[…]
Conceptually, $V(x)$ can be thought of as an energy function; points where $\nabla V(x_{\mathrm{data}})=0$ are near local minima of this energy, and the Hessian $\nabla^{2}V(x_{\mathrm{data}})$ provides information about local curvature (see Figure 4 for a qualitative illustration).
@@
[…]
@@ Appendix D Training details
Below, we detail the training configurations for CIFAR-10, ImageNet 32x32, CelebA, MNIST, and AAV. The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd [Ansel et al., 2024]. We optimize all models with the Adam optimizer Kingma and Ba [2014], and for CIFAR-10, ImageNet 32x32, and CelebA we maintain an exponential moving average (EMA) of the weights.
@@
$\beta$ $=\text{clamp threshold for $\mathcal{L}_{\mathrm{CD}}$ (i.e., we clamp $\mathcal{L}_{\mathrm{CD}}$ to be $\geq-\beta$).}$.
CIFAR-10.
@@
The architecture is shown in Figure 7. We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively time-independent) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output. Hyperparameters are: $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 200k iterations using Algorithm 1 with EMA $0.9999$ and then 10k more with Algorithm 2 and EMA $0.999$ on 4xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=8\times 10^{-3},\alpha=0.1,\beta=-0.05$.
ImageNet 32x32.
@@
The architecture is shown in Figure 7 (same as for CIFAR-10). Hyperparameters are: $\tau_{s}=1.5$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 600k iterations using Algorithm 1 with EMA $0.9999$ and then 7k more with Algorithm 2 and EMA $0.999$ on 7xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$ $\alpha=0.2,\beta=-0.01$.
CelebA.
@@
We scale the CIFAR-10 model by 2$\sim$ 2$\times$ 2 ×; see Figure 8. We keep $\tau_{s}=2.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 with EMA $0.9999$ then 7k with Algorithm 2 and EMA $0.999$ on 4xA100. The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
MNIST.
We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\Delta t=0.05$, $M_{\text{Langevin}}=50$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=0.25$. The architecture consists of a two-stage CNN followed by a two-layer MLP, comprising a total of 2M parameters.
AAV.
@@
We adopt the one-dimensional CNN architecture as used in [Bogensperger et al., 2025], summing the final-layer activations to obtain the potential. We train for 10k iterations using Algorithm 1 and for 1k iterations using Algorithm 2 on a single GeForce RTX 4090. The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$. For Algorithm 3 we use $\tau_{s}=1.7$ for AAV medium and $\tau_{s}=1.3$ for AAV hard, $\tau^{*}=0.9$, $\zeta=0.01$ for AAV medium and $\zeta=0.009$ for AAV hard and $\Delta t=0.01$. We set the target fitness to $y=1$ to aim for the maximum fitness in the generated sequences.
Input(3$\times$32$\times$32)UNet(37M params)Transformer(PatchEmbed+ViT, 12M params)ScalarPotential
@@
Figure 7: Diagram of our UNet+Transformer EBM for CIFAR-10 and ImageNet 32x32. A UNet (37M params) processes a 3$\times$32$\times$32 image; its output is fed into a Transformer head (PatchEmbed + 8-layer ViT, 12M params) that produces a scalar potential. Here we employ the identical UNet architecture as in [Tong et al., 2023], but with the time parameter fixed at $t=0$ to render the model time-independent.
Input(3$\times$64$\times$64)UNet(83M params)Transformer(PatchEmbed+ViT, 25M params)ScalarPotential
@@
Figure 8: Diagram of our UNet+Transformer EBM for CelebA. A UNet (83M params) processes a 3$\times$64$\times$64 image; its output is fed into a Transformer head (PatchEmbed+8-layer ViT, 25M params) that produces a scalar potential.

### v5

@@ A.1 Energy landscape during training
In Figure˜5, we visualize how the potential $V_{\theta}(x)$ transitions from a flow-like regime, where the OT loss enforces nearly zero curvature away from the data manifold (a), to an EBM-like regime, where the curvature around the new data geometry (here, two moons) is adaptively increased to approximate $\log p_{\text{data}}(x)$ (b). This two-stage design yields a well-shaped landscape that is both efficient to sample (thanks to a mostly flat potential between clusters) and accurate for density estimation near the data modes. For comparison, (c) shows an EBM trained solely with contrastive divergence, exhibiting sharper but less globally consistent basins.
@@
Figure 5: Visualization of the energy $V_{\theta}(x)$ landscapes driving the samples from eight Gaussians to two moons. See Figure˜1 for the 2D perspective. (a) The OT flow loss enforces zero curvature in $V_{\theta}(x)$ along the trajectories to the target. (b) Around the 2 Moons, the curvature of $V_{\theta}(x)$ is adjusted to approximate $\log p_{\text{moons}}(x)\propto V_{\theta}(x)$ while remaining close to the pretrained landscape elsewhere. Combining these objectives yields a potential energy landscape that is both efficient for sampling and representative of the underlying target data distribution. (c) An EBM is shown for comparison, trained using contrastive divergence loss. Visible mode collapse that slows down the equilibration. Less regular landscape away from the data as it needs many simulations to explore it.
@@ A.2 Ablation on the sampling time
Here, we provide ablation studies on CIFAR-10 unconditional generation. Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, producing a stable Boltzmann distribution from which one can sample. Figure˜6 illustrates the FID as a function of sampling time $\operatorname{\tau_{s}}$ for models trained under these different regimes. In the case of pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$; this occurs because, once the samples move close to the data manifold, there is no Boltzmann-like potential well to keep them from drifting away. Because the fidelity slope near the data manifold is steep with respect to sampling time, methods lacking explicit time-conditioning can easily overshoot or undershoot, significantly impacting fidelity. This behavior might explain why some models degrade in performance when made time-independent, as recently reported by Sun et al. [2025].
@@
In Figure˜6 we also report results for different values of the temperature-switching parameter $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation˜3).
1122330224466881010Sampling time $\operatorname{\tau_{s}}$FIDFID vs sampling time $\operatorname{\tau_{s}}$$\operatorname{\tau^{\ast}}=0.9,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\tau^{\ast}}=1.0,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
@@
Figure 6: CIFAR-10 unconditional generation FID vs. sampling time $\operatorname{\tau_{s}}$ when sampling from models trained under different scenarios: pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ and combined $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, with temperature regime switching parameter $\operatorname{\tau^{\ast}}\in\{0.9,1.0\}$ during sampling. Lower FID indicates better generative quality. All results for Euler-Heun with $\Delta t=0.01$.
A.3 Ablations on the OT Solver
We evaluate the computational overhead and sensitivity to solver choice of the OT solver employed in our experiments.
Computational overhead.
In the CIFAR-10 experiment, the OT solver accounts for roughly 1.5% of the training iteration time during Phase 1 Algorithm˜1. This overhead decreases to a negligible level (approximately 0.01%) in Phase 2 Algorithm˜2, where computational costs are predominantly dominated by the generation of negative samples.
Impact of solver accuracy and complexity.
Let us (over‑idealise) and model a standardised CIFAR‑10 image as a vector $x\in\mathbb{R}^{d}$, drawn from $\mathcal{N}(0,I_{d})$, with dimensionality $d=32\times 32\times 3=3072$, and paired with an independent Gaussian-noise vector $z\sim\mathcal{N}(0,I_{d})$. Each coordinate of the difference (zx)(z-x) thus follows $\mathcal{N}(0,2)$, and the squared Euclidean distance distribution is $|z-x|_{2}^{2}\sim 2\chi_{d}^{2}$, which has mean $2d$, standard deviation $\sqrt{8d}$, and relative spread $\frac{\sqrt{8d}}{2d}\approx 0.025$. This demonstrates the "thin-shell" phenomenon, implying that all entries of the cost matrix $C_{ij}=|z_{i}-x_{j}|^{2}_{2}$ concentrate around nearly identical values, consistent with the distance-concentration effect observed by [Aggarwal et al., 2001]. Consequently, the choice among exact linear programming (LP), entropic regularisation (Sinkhorn), or even random matching should yield nearly identical cumulative optimal-transport costs, despite their different complexities: $O(n^{3}\log n)$ for LP, $O(\frac{n^{2}}{\kappa})$ for Sinkhorn (with regularisation strength $\kappa$), and $O(n)$ for random matching.
Empirical evidence supporting this is summarised below:
  • Our CIFAR-10 runs: FID degrades from 3.34 (LP) to 3.37 (random).
  • [Tong et al., 2023] Table 5: FID scores of 4.44 (LP) vs 4.46 (random) at 100 steps.
  • [Tong et al., 2023] Fig. D.2: solver accuracy saturates beyond batch size 16 in the 2D task.
  • [Terpin et al., 2024] App. C.2: LP and Sinkhorn methods produce indistinguishable results unless regularisation is extreme.
We employ the LP approach for robustness, negligible cost, and no additional hyperparameters.
A.4 Sampling Time Analysis
Computing the gradient $\nabla_{x}V_{\theta}(x)$ via automatic differentiation (autograd [Ansel et al., 2024]) introduces additional computational overhead compared to directly evaluating $V_{\theta}(x)$. Specifically, on the CIFAR-10 network architecture (see Figure˜7), gradient evaluation is approximately $2.15\times$ slower, as it requires both forward and backward passes. In contrast, flow matching and diffusion models directly parameterize the velocity field, thus only requiring forward computations during sampling.
Nevertheless, despite this per-step computational cost, our method achieves competitive overall sampling efficiency due to a reduced number of integration steps needed for high-quality generation. As demonstrated in Table˜4, Energy Matching achieves a lower FID (3.34) in 173 seconds per batch, outperforming OT-FM (FID 3.74 in 136 seconds per batch) and DDPM++ (FID 3.45 in 183 seconds per batch). Our results thus indicate a favorable balance between computational overhead per step and total sampling runtime.
Table 4: Comparison of sampling efficiency and quality on CIFAR-10 (batch size 128, NVIDIA R6000 48GB GPU). Despite gradient computation overhead ($\nabla_{x}V_{\theta}(x)$ via backward pass), Energy Matching achieves superior FID scores with competitive wall-clock sampling time.
Method Params Steps Sampling Time [s]↓ FID↓
Flow-/Diffusion-based Models
     OT-FM [Tong et al., 2023] 37M 1000 136 3.74
     DDPM++ [Kim et al., 2021] 62M 1000 183 3.45
Energy-based Models
     Energy Matching (Ours) 50M 325 173 3.34
Appendix B Details on AAV inverse design protein generation
@@
We optimize protein fitness for adeno-associated virus (AAV) sequences in the medium and hard data regimes proposed by [Kirjner et al., 2024], using latent encodings and backbone architectures from [Bogensperger et al., 2025]. Conditional sampling employs classifier guidance via learned predictor networks $g_{\phi}$ or $\tilde{g}_{\phi}$ to steer samples toward high-fitness regions. The CNN-based fitness predictors from [Kirjner et al., 2024] are trained only on the limited training data for each regime.
@@
Training follows Algorithm˜1 and Algorithm˜2. We sample 128 sequences using Algorithm˜3, keeping the same batch size across all baselines; detailed hyperparameters are given in Appendix˜D. Generated sequences are evaluated for fitness using the learned oracle from [Kirjner et al., 2024], and further assessed for both intra-set diversity and novelty relative to the training sequences [Jain et al., 2022]. Our approach achieves state-of-the-art fitness while improving diversity (see Table˜5). Incorporating interaction energy in Algorithm˜3 further enhances diversity with manageable impact on fitness.
@@ Appendix C Details on LID estimation
$\mathrm{LID}(x_{\mathrm{data}})=d-\lim_{r\to 0}\frac{\log\bigl(M(r)\bigr.)}{\log(r)}.$
@@
$\nabla V(x_{\mathrm{data}})\approx 0\quad\text{and}\quad\nabla^{2}V(x_{\mathrm{data}})\text{ is positive semidefinite (or nearly so).}$
Conceptually, $V(x)$ can be thought of as an energy function; points where $\nabla V(x_{\mathrm{data}})=0$ are near local minima of this energy, and the Hessian $\nabla^{2}V(x_{\mathrm{data}})$ provides information about local curvature (see Figure˜4 for a qualitative illustration).
@@
$M(r)=\int_{B(x_{\text{data}},r)}p(x)dx\propto\int_{B(x_{\text{data}},r)}\exp\left(-\tfrac{1}{2\varepsilon}(x-x_{\mathrm{data}})^{\top}\nabla_{x}^{2}V(x_{\mathrm{data}})(x-x_{\mathrm{data}})\right)dx.$
@@ Appendix D Training details
Below, we detail the training configurations for CIFAR-10, ImageNet 32x32, CelebA, MNIST, and AAV. Additionally, we provide intuitions for practical hyperparameter choices to facilitate effective training across additional datasets. We recommend using SiLU activation functions wherever possible, as they smooth out the energy landscape and improve the numerical stability of the $\nabla_{x}V(x)$ computation. The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd [Ansel et al., 2024]. We optimize all models using the Adam optimizer [Kingma and Ba, 2014] and maintain an exponential moving average (EMA) of the model weights.
@@
[…]
CIFAR-10:
@@
The architecture is shown in Figure˜7. We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively time-independent) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output. Hyperparameters are: $\tau_{s}=3.25$, $\tau^{*}=1.0$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 145k iterations using Algorithm˜1 with EMA $0.9999$ and then 2k more with Algorithm˜2 and EMA $0.99$ on 4xA100. The batch size is 128, learning rate is $1.2\times 10^{-3}$, $\varepsilon_{\max}=0.01$, $\lambda_{\mathrm{CD}}=1\times 10^{-3}$, $\alpha=0.1$, and $\beta=0.02$. Negatives initialized on the data manifold follow the same temperature schedule as those initialized from the noise.
ImageNet 32x32:
@@
The architecture is shown in Figure˜7 (same as for CIFAR-10). Hyperparameters are: $\tau_{s}=2.5$, $\tau^{*}=1.0$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 640k iterations using Algorithm˜1 with EMA $0.9999$ and then 1k more with Algorithm˜2 and EMA $0.99$ on 7xA100. The batch size is 128, learning rate is $6\times 10^{-4}$, $\varepsilon_{\max}=0.01$, $\lambda_{\mathrm{CD}}=1\times 10^{-3}$, $\alpha=0.1$, and $\beta=0.02$. Negatives initialized on the data manifold follow the same temperature schedule as those initialized from the noise.
CelebA:
@@
We scale the CIFAR-10 model by $\sim 2\times$; see Figure˜8. We set $\tau_{s}=2.0$, $\tau^{*}=1.0$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm˜1 with EMA $0.9999$ then 4k with Algorithm˜2 and EMA $0.99$ on 4xA100. The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
MNIST:
We downscale the CIFAR-10 model (Figure˜7) to 2M parameters by reducing the UNet base width to 32 channels, using channel multipliers [1, 2, 2], setting the number of attention heads in the UNet to 2, simplifying the Transformer head to an embedding dimension of 128, 2 Transformer layers, 2 attention heads, and adjusting the output scale to 100.0. We set $\tau_{s}=2.0$, $\tau^{*}=1.0$, $\Delta t=0.025$, $M_{\text{Langevin}}=75$, and train for 50k iterations using Algorithm˜1 with EMA $0.999$ then 3.3k with Algorithm˜2 and EMA $0.99$ on a single A100. The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, $\lambda_{\mathrm{CD}}=1\times 10^{-3}$, $\alpha=0.0$, and $\beta=0.05$. Negatives initialized on the data manifold follow the same temperature schedule as those initialized from the noise.
AAV:
@@
We adopt the one-dimensional CNN architecture as used in [Bogensperger et al., 2025], summing the final-layer activations to obtain the potential. We train for 10k iterations using Algorithm˜1 and for 1k iterations using Algorithm˜2 on a single A100. The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, $M_{\text{Langevin}}=200$, $\Delta t=0.01$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$. For Algorithm˜3 we use $\tau_{s}=1.7$ for AAV medium and $\tau_{s}=1.3$ for AAV hard, $\tau^{*}=0.9$, $\zeta=0.01$ for AAV medium and $\zeta=0.009$ for AAV hard. We set the target fitness to $y=1$ to aim for the maximum fitness in the generated sequences.
Intuition for other datasets:
It is essential for negative samples to reach the equilibrium distribution induced by the model or at least the proximity of the data manifold. The condition $M_{\text{Langevin}}\times\Delta t\gg 1$ is critical to achieving this, with $\Delta t$ small enough to ensure that negative samples remain of sufficient quality—typically the same $\Delta t$ as used in flow matching for the given generation task. In practice, we set $M_{\text{Langevin}}\times\Delta t=2$ across most experiments. We recommend starting with $\tau^{*}=1.0$ to ensure optimal transport regularization near the data manifold, thereby enhancing training stability. If additional conditions are required during sampling, exploring lower values ($\tau^{*}<1.0$) may be beneficial, as this parameter does not need to remain consistent between training and sampling (as shown in Figure˜6). Training with $\tau^{*}<1.0$ is advised only in special cases, such as extremely low-dimensional problems like that shown in Figure˜1, where it is possible to simultaneously be far from the data manifold (from the perspective of the target mode) and close to it (from the perspective of another mode). The parameter $\varepsilon_{\max}$ controls how extensively negative samples explore the space. For unconditional generation, we use $\varepsilon_{\max}=0.01$, but for inverse problems or design tasks, higher values (e.g., $\varepsilon_{\max}=0.05$) can improve robustness. The parameter $\tau_{s}$ significantly depends on the task (unconditional or conditional) and thus must be tuned accordingly post-training. Without explicit tuning, selecting $\tau_{s}=M_{\text{Langevin}}\times\Delta t$ is a reasonable default. Finally, parameters $\lambda_{\mathrm{CD}}$, $\alpha$, and $\beta$ influence the stability of contrastive training and must be empirically determined alongside appropriate early stopping.
[…]
@@
[…]
@@
[…]